find the sum of the degree of the measures of the interior angles of a regular polygon that has 17 sides
Answer by jim_thompson5910(35256)

S = 180(n-2)

S = 180(17-2)

S = 180(15)

S = 2700


So the sum of all of the measures of the interior angles of a 17-sided polygon is 2700 degrees.